1. Determine the lateral area of a cylinder of height 300mm and base radius of 100mm.

2. A right circular cylinder with an altitude of 40 cm and base diameter of 30cm has a right circular cone at its top with the same base. If the cone is 60cm high, determine its lateral area.
1
Expert's answer
2020-12-15T07:16:28-0500

"h = 300\; mm \\

r = 100 \;mm \\

Lateral \;area\; of \;cylinder = 2πrh \\

= 2 \times 3.14 \times 100 \times 300 \\

= 188520 \; mm^2"

Answer: 188520 mm2

The lateral surface area "S_1" of a right circular cylinder with an altitude of "h_1=40\ cm" and base diameter of "d=30\ cm" is equal to "S_1=\pi \cdot d\cdot h=\pi\cdot 40\cdot 30=1200\pi" ("cm^2" )

The lateral surface area "S_2" of a right circular cone with an altitude of "h_2=60\ cm" and base diameter of "d=30\ cm" is equal to "S_2=\pi \cdot r\cdot l", where "r=\frac{d}{2}=15\ (cm)" is the radius of the base and "l=\sqrt{(h_2)^2+r^2}=\sqrt{60^2+15^2}=\sqrt{3600+225}=\sqrt{3825}=15\sqrt{17}\ (cm)" is the slant height of the cone Therefore, "S_2=\pi\cdot 15\cdot 15\sqrt{17}=225\sqrt{17}\pi\ (cm^2)".

The lateral surface area of a right circular cylinder with a right circular cone at its top is"1200\pi+225\sqrt{17}\pi=75\pi(16+3\sqrt{17})\ (cm^2)."